# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

### v2

@@ Introduction
12:   else  Train model on $\omega^{*}$. $N(\omega_{p}):=N(\omega_{p})+1$.
@@
Implementation Details. Our SAMPL method is built upon RLIPv2 (Yuan et al. 2023) with a Swin-Tiny (Liu et al. 2021) backbone and has $214M$ trainable parameters. The backbone is pretrained on Visual Genome (Krishna et al. 2017), MS-COCO (Lin et al. 2014) and Object365 (Shao et al. 2019), and fine-tuned on the HICO-DET or V-COCO datasets. We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps. Models are trained for $80$ epochs using cross-entropy loss for object classification, focal loss for verb classification, and L1 + GIoU loss for bounding, with KL divergence for latent space regularization. The encoder and decoder consist of $6$ and $3$ layers, respectively. For the dynamic severity adjustment mechanism, we set the initial threshold as $\tau_{\text{init}}=0.15$ and $\epsilon=10^{-6}$.
@@
Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks. The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions. Corruptions are denoted as Abbreviation, i.e., $L\gamma$, where $\gamma$ represents the severity level. Details are shown in Sec. Benchmark.
@@
Analysis of the Benchmark. Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET (Wang et al. 2020) and V-COCO (Gupta and Malik 2015) datasets. SAMPL surpasses RLIPv2 (Yuan et al. 2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations. On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy. These results are further supported by Fig. 2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary). While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods by progressively training the model with semantic-aware occlusions, enabling it to learn context-driven features that are resilient to diverse and severe corruptions. These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning. Through a comprehensive analysis of representative HOI models, we derive the following key findings.
Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models. Experiments on the clean and corrupted test set (Table 1) reveal that transformer-based architectures exhibit robustness variations dependent on their pretraining strategies. RLIPv2, which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT, SOV-STG and other transformer-based methods. This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks. Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT, show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
@@
Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 (Yuan et al. 2023). The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images. Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
Observation 2: Object localization strategies dictate robustness in two-stage architectures. Two-stage methods show varying robustness across different object localization strategies. Region-proposal-based models (Zhang et al. 2021) achieve high clean-set accuracy but degrade under corruption due to error propagation in localization. In contrast, query-based methods (Zhang, Campbell, and Gould 2022) adapt object representations dynamically, enhancing robustness on HICO-DET. However, on V-COCO, UPT with a larger backbone surprisingly shows lower CRI than CDN with a smaller backbone, likely attributed to V-COCO’s limited interaction diversity. This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
@@
Table 2: Ablation study on various perturbation strategies during training. We evaluate SAM-guided HOI-aware Semantic Masking (SAMSM), Adaptive Gaussian Noise Perturbation (AGNP), and Randomized Pixelation Perturbation (RPP) on the V-COCO.
Observation 3: Interaction representation entanglement amplifies robustness degradation under perturbations. Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation. MUREN, with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness. Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s adaptive query-based approach improves robustness by dynamically refining predictions. This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
Observation 4: Dataset interaction diversity directly influences robustness generalization. Models trained on HIO-DET achieve higher MRI and CRI than V-COCO-trained models due to greater interaction diversity, enhancing resilience. However, SAMPL shows stronger robustness gains on V-COCO, as its progressive learning strategy compensates for dataset limitations, achieving comparable robustness with less data and fewer iterations. This underscores the importance of diverse interactions for general robustness and adaptive training for faster improvements in smaller datasets. Structural design considerations in both architectures and training paradigms are crucial for balancing recognition accuracy and robustness in real-world applications.
@@
Table 3: Ablation study on training with different severity level combinations. SAMSM masks each instance based on 40%–80% of its bounding box height and width, using three candidate levels during training.
Analysis of the Ablation Study. To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset. The results, presented in Tables 2 and 3, analyze the effects of different perturbation types and perturbation level settings on model performance.
Ablation of the Semantic-Aware Masking: Table 2 evaluates three perturbation strategies applied independently during training, excluding our progressive learning method to isolate their direct impact. SAM-guided HOI aware Semantic Masking (SAMSM) is applied in 50% of training epochs, occluding instance’s regions while preserving contextual cues. Adaptive Gaussian Noise Perturbation (AGNP) introduces feature-level noise with a 50% probability, while Randomized Pixelation Perturbation (RPP) pixelates input images with the same probability as data augmentation. SAMSM achieves the highest MRI and CRI scores, suggesting that structured occlusions enhance robustness by enforcing relational reasoning under missing object regions. Unlike AGNP, which perturbs latent features without spatial structure, SAMSM provides localized challenges that align with real-world occlusions. Despite expectations that RPP would improve robustness—given that pixelation causes the most significant performance drop in baseline corruption tests (as shown in Fig. 4)—its randomized nature indiscriminately distorts both essential and redundant information, making adaptation less effective. These results suggest that robustness in HOI detection benefits more from structured perturbations that challenge object-awareness rather than indiscriminate degradations that introduce confounding artifacts.
Figure 4: Performance degradation of RLIPv2 (Yuan et al. 2023) on HICO-DET and V-COCO under RoHOI. The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO. Clean dataset performance serves as a reference, with shaded areas indicating corruption impact.
Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness. SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores. Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($70\%$) disrupts spatial structures, reducing discriminative capacity. These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
@@
Conclusion
In this paper, we introduced RoHOI, the first robustness benchmark for HOI detection, evaluating models under $20$ real-world perturbations. Our analysis shows that existing models suffer significant degradation, highlighting the need for more robust approaches. To address this, we propose SAMPL, a Semantic-Aware Masking-based Progressive Learning method that utilizes SAM-generated region-aware masks and dynamically adjusts masking severity. SAMPL enhances robustness while maintaining high accuracy on clean datasets, outperforming state-of-the-art methods. By bridging the gap between clean-data performance and real-world robustness, our work advances reliable and adaptable HOI detection models for practical use.
Appendix A Social Impact and Limitations
@@
Appendix B Standard HOI Evaluation Metrics
This section provides a detailed explanation of the evaluation metrics used for HICO-DET (Xu et al. 2019) and V-COCO (Gupta and Malik 2015) in our main paper. These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
$\bm{\mathrm{mAP}}\,\textbf{(Full, Rare, Non-Rare)}$
@@ Appendix B Standard HOI Evaluation Metrics
$\bm{\mathrm{AP}^{\#1}_{\text{role}}}$ and $\bm{\mathrm{AP}^{\#2}_{\text{role}}}$
@@
Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
CDN-L (Zhang et al. 2021) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
GEN-VLKT-L (Liao et al. 2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
MUREN (Kim, Jung, and Cho 2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
RLIPv2-ParSeDA (Yuan et al. 2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
[…]
@@
Appendix C Corruption Types in the Dataset
Optical System-Induced (OS) Artifacts
@@ Appendix C Corruption Types in the Dataset
Sensor, Compression and Transmission (SCT) Artifacts
@@
Environmentally Induced (EI) Artifacts
@@
Geometric and Scene (G&S) Distortions
@@
Appendix D Benchmark Results on Corruption Types

### v3

@@ 4 Method
12:   else  Train model on $\omega^{*}$. $N(\omega^{*}):=N(\omega^{*})+1$.
@@
Table 1: Comparison of HOI detection methods on HICO-DET and V-COCO under clean and corrupted settings. Results on HICO-DET are reported on Full, Rare and Non-Rare splits; V-COCO results use $\mathrm{AP}^{\#1}_{\text{role}}$ and $\mathrm{AP}^{\#2}_{\text{role}}$. Columns Backb., Det., VLM indicate whether a model uses a pretrained backbone, a detection-model pretraining, or vision–language pretraining. MRI and CRI quantify robustness under corruptions (higher is better).
@@ 5 Experimental Results
Implementation Details. Our SAMPL method is built upon RLIPv2 Yuan et al. (2023) with a Swin-Tiny Liu et al. (2021) backbone and has $214M$ trainable parameters. The backbone is pretrained on Visual Genome Krishna et al. (2017), MS-COCO Lin et al. (2014) and Object365 Shao et al. (2019), and fine-tuned on the HICO-DET or V-COCO datasets. We use the AdamW optimizer Loshchilov & Hutter (2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps. Models are trained for $80$ epochs using cross-entropy loss for object classification, focal loss for verb classification, and L1 + GIoU loss for bounding, with KL divergence for latent space regularization. The encoder and decoder consist of $6$ and $3$ layers, respectively. For the dynamic severity adjustment mechanism, we set the initial threshold as $\tau_{\text{init}}=0.15$ and $\epsilon=10^{-6}$.
(a) Refer to caption
(b) Refer to caption
@@
Figure 2: (a) Performance degradation of RLIPv2 Yuan et al. (2023) on HICO-DET and V-COCO under RoHOI. The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO. Clean dataset performance serves as a reference, with shaded areas indicating corruption impact. (b) Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks. The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions. Corruptions are denoted as Abbreviation, i.e., $L\gamma$, where $\gamma$ represents the severity level. Details are shown in Sec. 3 and Appendix.
@@
Analysis of the Benchmark. Tab. 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET Wang et al. (2020) and V-COCO Gupta & Malik (2015) datasets. SAMPL surpasses RLIPv2 Yuan et al. (2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations. On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy. These results are further supported by Fig. 2 (b), where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the Appendix). While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods by progressively training the model with semantic-aware occlusions, enabling it to learn context-driven features that are resilient to diverse and severe corruptions. These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning. Through a comprehensive analysis of representative HOI models, we derive the following key findings.
Observation 1: Vision–language relational pretraining is associated with higher corruption-robustness. Within each dataset (Tab. 1), HOI detectors trained with relational vision–language objectives occupy the top tier of MRI/CRI, while transformer-based counterparts without such cross-modal relational supervision lag behind. A plausible mechanism is that relational fusion during pretraining aligns verb–object evidence and regularizes interaction features against distribution shift; We interpret this pattern as associative rather than causal; backbone capacity remains a potential confounder.
@@
Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 Yuan et al. (2023). The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images. Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
[…]
Observation 4: Progressive semantic masking primarily improves stability (CRI) and smooths severity–performance curves. Relative to a strong vision–language pretrained baseline in Tab. 1, SAMPL consistently increases CRI within each dataset, while MRI gains are moderate or dataset-dependent. This pattern aligns with evidence that self-attentive encoders are sensitive to patch-level corruptions: structure-aware, progressive masking encourages attention to retain relational evidence under partial visibility, yielding smoother severity–performance profiles without sacrificing clean accuracy.
Table 2: Ablation studies on V-COCO. (a) Ablation on various perturbation strategies during training: SAM-guided HOI-aware Semantic Masking (SAMSM), Adaptive Gaussian Noise Perturbation (AGNP), and Randomized Pixelation Perturbation (RPP). (b) Ablation on training with different severity level combinations for SAMSM: each instance is masked based on 40%–80% of its bounding box height and width, using three candidate levels during training.
(a) Training perturbation strategies
@@
(b) SAMSM cover ratio combinations
@@
Analysis of the Ablation Study. To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset. The results, presented in Tab. 2, analyze the effects of different perturbation types and perturbation level settings on model performance.
Ablation of the Semantic-Aware Masking: Tab. 2 (a) evaluates three perturbation strategies applied independently during training, excluding our progressive learning method to isolate their direct impact. SAM guided HOI-aware Semantic Masking (SAMSM) is applied in 50% of training epochs, occluding instances’ regions while preserving contextual cues. Adaptive Gaussian Noise Perturbation (AGNP) introduces feature-level noise with a 50% probability, while Randomized Pixelation Perturbation (RPP) pixelates input images with the same probability as data augmentation. SAMSM achieves the highest MRI and CRI scores, suggesting that structured occlusions enhance robustness by enforcing relational reasoning under missing object regions. Unlike AGNP, which perturbs latent features without spatial structure, SAMSM provides localized challenges that align with real-world occlusions. Despite expectations that RPP would improve robustness—given that pixelation causes the most significant performance drop in baseline corruption tests (as shown in Fig. 2 (a))—its randomized nature indiscriminately distorts both essential and redundant information, making adaptation less effective. These results suggest that robustness in HOI detection benefits more from structured perturbations that challenge object-awareness rather than indiscriminate degradations that introduce confounding artifacts.
Effect of Perturbation Levels: Tab. 2 (b) analyzes the impact of masking severity on robustness. SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores. Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($70\%$) disrupts spatial structures, reducing discriminative capacity. These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
@@
6 Conclusion
In this work, we introduced RoHOI, the first benchmark dedicated to evaluating the robustness of Human-Object Interaction (HOI) detection models under 20 diverse real-world corruptions. Our analysis revealed that even state-of-the-art approaches degrade substantially under perturbations, underscoring the need for robustness-focused evaluation. To address this challenge, we proposed SAMPL (Semantic-Aware Masking-based Progressive Learning), which leverages SAM-generated region-aware masks and a dynamic progressive masking strategy to strengthen resilience. Extensive experiments on HICO-DET and V-COCO showed that SAMPL consistently outperforms strong baselines under corruption while preserving state-of-the-art performance on clean benchmarks. Beyond performance gains, our results highlight that clean accuracy and robustness are only partially coupled, that structured perturbations provide more effective robustness cues than random noise, and that architecture design plays a crucial role in stability. While RoHOI captures a broad range of corruptions, it still focuses on synthetic perturbations; future work should explore robustness under real-world domain shifts, extreme degradations, and fairness-aware evaluation. We envision RoHOI as a catalyst for advancing robustness research in HOI detection, and SAMPL as an initial step toward building reliable and deployable models for safety-critical applications.
Acknowledgements
The project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1574 – 471687386. This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026), the MWK through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, and in part by the BMBF through a fellowship within the IFI program of the German Academic Exchange Service (DAAD), in part by the HoreKA@KIT supercomputer partition, in part by the National Natural Science Foundation of China (No. 62473139), in part by the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20).
Appendix A Appendix
A.1 Social Impact and Limitations
@@
A.2 Standard HOI Evaluation Metrics
This section provides a detailed explanation of the evaluation metrics used for HICO-DET Xu et al. (2019) and V-COCO Gupta & Malik (2015) in our main paper. These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
A.2.1 $\mathbf{mAP}$: Full, Rare, Non–Rare
@@
A.2.2 $\mathbf{AP}^{\#1}_{\text{role}}\ \text{and}\ \mathbf{AP}^{\#2}_{\text{role}}$
@@
A.3 Corruption Types in the Dataset
A.3.1 Optical System-Induced (OS) Artifacts
@@
A.3.2 Sensor, Compression and Transmission (SCT) Artifacts
@@
A.3.3 Environmentally Induced (EI) Artifacts
@@
A.3.4 Geometric and Scene (G&S) Distortions
@@ A.4 Purpose of Inclusion
A.5 Benchmark Results on Corruption Types
@@
Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
CDN-L Zhang et al. (2021a) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
GEN-VLKT-L Liao et al. (2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
MUREN Kim et al. (2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
RLIPv2-ParSeDA Yuan et al. (2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
[…]